# Where to Begin: Efficient Pretraining via Subnetwork Selection and Distillation

### v1

@@ Abstract
Small Language models (SLMs) offer an efficient and accessible alternative to Large Language Models (LLMs), delivering strong performance while using far fewer resources. We introduce a simple and effective framework for pretraining SLMs that brings together three complementary ideas. First, we identify structurally sparse sub-network initializations that consistently outperform randomly initialized models of similar size under the same compute budget. Second, we use evolutionary search to automatically discover high-quality sub-network initializations, providing better starting points for pretraining. Third, we apply knowledge distillation from larger teacher models to speed up training and improve generalization. Together, these components make SLM pretraining substantially more efficient: our best model, discovered using evolutionary search and initialized with LLM weights, matches the validation perplexity of a comparable Pythia SLM while requiring $\mathbf{9.2\times}$ fewer pretraining tokens. We release all code and models at https://github.com/whittle-org/whittle/, offering a practical and reproducible path toward cost-efficient small language model development at scale.
@@ 1 Introduction
  • Comprehensive analysis. We provide the first systematic comparison of sub-network initialization under knowledge distillation versus standard cross-entropy training. Our study spans multiple student scales and investigates how teacher size influences effectiveness.
  • Reproducible framework. We release an open-source library [^1] https://github.com/whittle-org/whittle/ for extracting sub-networks from existing LLM checkpoints. Together with our empirical findings, this establishes practical guidelines for compute-optimal SLM pretraining across different scales.
@@ 2.3 SLM Pretraining and Distillation
$\mathcal{L}_{\text{D}}=\sum_{i}p_{t_{i}}^{(T)}\log\frac{p_{t_{i}}^{(T)}}{p_{s_{i}}^{(T)}},$ (3)
@@ 3 Whittle: A Library for SLM Pre-training and Distillation
Given a pretrained decoder-only LLM from litgpt, we first convert it into a whittle model to enable flexible sub-network extraction. To evaluate a sub-network using the whittle model, we dynamically activate only structured components of the LLM associated with that sub-network using the set_sub_network() API (Listing 1). It allows the user to explicitly set architectural parameters of the sub-network, such as embedding dimension, intermediate size, number of heads, layers, query groups, and head size, as well as indices for sampled neurons, layers, and heads. Importantly, it allows to vary the number of heads, head size, intermediate size, and query groups across layers. This function is a core utility in whittle, supporting downstream procedures such as search, pretraining, and distillation.
@@
The search() API (Listing 3, Appendix F) constructs a whittle super-network from a base HuggingFace model and facilitates automated sub-network selection. It supports evolutionary strategies as well as algorithms from syne-tune (Salinas et al., 2022) [^4] https://github.com/syne-tune/syne-tune, and performs constrained search across parameter bins via rejection sampling. Each candidate sub-network is instantiated through set_sub_network() and evaluated on a task-specific metric, such as perplexity, to guide the search process.
@@
The convert_subnet_to_litgpt_model() function (Listing 4, Appendix F) transforms a selected sub-network configuration into a standalone GPT model within the litgpt framework. Given a super-network and a dictionary specifying architectural configurations (e.g., embedding dimension and number of heads), this utility extracts the corresponding sub-network and instantiates it as an independent GPT model. The resulting model can then be employed for downstream tasks such as pretraining, fine-tuning, or distillation.
@@
The pretrain() function (Listing 2, Appendix F) enables pretraining of a sub-network initialized from a checkpointed GPT model. Given the model weights, a configuration file describing the sub-network architecture, and a target dataset, this utility restores the model and resumes training from the specified state.
@@ 4.1 Search Space Definitions
Table 1 summarizes the search spaces for Pythia-6.9B and Pythia-12B, listing the allowable values for each transformer dimension. In the coarse layer-wise and fine-grained layer-wise settings, $h$, $h_{s}$, and $d$ are sampled independently at each layer. The fine-grained spaces extend this further with neuron-level sampling within each dimension, as detailed in Section 2.1.
@@ 4.2 Evolutionary Search for Optimal SLMs
We use perplexity on wikitext (Merity et al., 2017) as selection metric, with mutation and crossover probabilities fixed at $0.2$. For Pythia-6.9B and 12B, we define three bins with parameter counts of 385M–426M (bin-1), 961M–1.06B (bin-2), and 2.64B–2.91B (bin-3), centered at 5% of Pythia-410M, Pythia-1B, and Pythia-2.8B, respectively [^7] The bins were computed based on the exact number of parameters in the Pythia models. Each evolutionary run proceeds for 100 epochs, with results in the fine-grained layer-wise setting reported at the final epoch before rejection sampling becomes infeasible due to the combinatorial growth of the search space.
@@ 4.3 Pretraining of SLMs
Figure 4 presents the pretraining results. We compare pretraining of the extracted best sub-network (Supernet-init) against two baselines: (i) Random-init, where the same architecture is trained with random initialization and a 10B-token budget, and (ii) the original Pythia model (center of the bin), also trained with random initialization and the same budget. Across parameter bins, initializing from the supernet yields consistent improvements in validation perplexity. Notably, the gains are most pronounced for bin-3, indicating that supernet initialization is particularly beneficial in higher-parameter regimes, where our model achieved the same validation perplexity with $9.2\times$ fewer total training tokens.
@@ 4.4 Distillation of SLMs
Figure 5 illustrates the effect of distillation. We find that distillation consistently improves perplexity in both bin-1 and bin-2, with the model in bin-2 achieving the same performance as Pythia-1B with $5.11\times$ fewer tokens.
@@ 5 Ablations
Finally, in Figure 8, we evaluate different search metrics. Specifically, we compare activation-based importance scores (as in Minitron (Muralidharan et al., 2024)) and weight-magnitude scores (Han et al., 2015) against directly optimizing for perplexity in our setup. We define the details of the importance score computation procedure, i.e. the metric guiding the search, in Appendix E. All searches are run with the same 16-hour budget per parameter bin. We find that perplexity-based search consistently achieves lower perplexity than proxy metrics, suggesting that importance and magnitude scores are less reliable indicators of sub-network quality.
@@ Appendix A Related Work
Appendix B Hyperparameter Configurations of Experiments
@@ Appendix D Additional Results
Teacher selection.
We also examine how the size of the teacher impacts student efficiency under the same compute budget (Figure 12). The smaller teacher model (Pythia-1B) achieves end-of-training perplexity comparable to, and marginally lower than, that of the larger Pythia-6.9B: $35.06$ vs. $35.15$ at 22B tokens. This indicates that, for our setup, alignment of inductive biases and gradient signal quality may be more critical than raw size of the teacher, and that excessively large teachers may be unnecessary under budget-constrained distillation.
@@
Table 10: Evaluation of sub-networks extracted from Pythia-6.9B for bin-1 and bin-2. Reported numbers are metrics as defined in Section 4.5 (%). We compare training with the cross entropy loss ( from-supernet) to training with knowledge distillation (from-supernet-distill) loss.
@@ Appendix E Details on Importance Scoring
Importance scoring aims at defining scores for each transformer dimension, neuron or architecture parameter based on activation or weight magnitude. In our case, for a sub-network, the corresponding importance score serves as the proxy to sub-network quality or performance metrics like perplexity. Specifically, higher the importance score of a sub-network, superior the quality of the sub-network.
We adopt the dimension-wise importance scoring proposed by Muralidharan et al. [2024], which uses the activation of a component as proxy for its importance.
Given a batch as input ${\bm{X}}\in\mathbb{R}^{B\times T\times d_{model}}$ after applying the embedding layer ${\bm{W}}^{emb}$ we compute the following scores for each component, where $B$ corresponds to the batch dimension and $T$ corresponds to the sequence length dimension, and absabs corresponds to the absolute value function:
@@ Appendix F Whittle APIs
In addition to the functionalities core to our framework as defined in Section 3, we additionally define an api to compute different importance metrics for different sub-network dimensions. An overview of the Whittle library is shown in Figure 9.
@@
Below we present the details of our API design for setting subnetwork 1, pretrain 2, convert to litgpt 4, distillation 5, 3 search and importance metric computation 6.
Listing 1: API for activating a sub-network
def set_sub_network(
sub_network_n_embd: int = 4, # Embedding dim
sub_network_intermediate_size: int | list[int] = 42 # MLP size
sub_network_num_heads: int | list[int] = 4, # No. of Attention heads
sub_network_n_layers: int = 4, # No. of Layers
sub_network_query_groups: int | list[int] = 2, # No. of Query groups
sub_network_head_size: int | list[int] = 4, # Head size
sampled_layer_indices: list[int] = [2,3,5,6], # Sampled layers
Listing 2: API for pretraining.
@@
Listing 3: API for supernetwork search.
@@
Listing 4: API for converting a subnet into a LitGPT model.
@@
Appendix G Reproducibility Statements
@@
  • Code and Repository: We release all our code and scripts to reproduce our experiments at https://github.com/whittle-org/whittle/
  • Datasets and Pretrained Models: We evaluate on available benchmarks from lm-eval-harness (https://github.com/EleutherAI/lm-evaluation-harness) and use the publicly available Nemotron-CC dataset https://research.nvidia.com/labs/adlr/Nemotron-CC/ for training. Furthermore we use the Pythia-model suite, which is open-source https://github.com/EleutherAI/pythia.
  • Compute Resources: All our search experiments were run on on L40 GPU per parameter bin and base model. All our pretraining runs for bin-1 and bin-2 were run on 8 L40 GPUs and bin-3 was run on 4 H200 GPUs. All our distillation experiments were run on 4 H200 GPUs. We use cuda version 11.8.
  • Evaluation and Artifacts: Upon acceptance of the paper we will publicly release model checkpoints for all our experiments.

### v2

@@ Abstract
Small Language models (SLMs) offer an efficient and accessible alternative to Large Language Models (LLMs), delivering strong performance while using far fewer resources. We introduce a simple and effective framework for pretraining SLMs that brings together three complementary ideas. First, we identify structurally sparse sub-network initializations that consistently outperform randomly initialized models of similar size under the same compute budget. Second, we use evolutionary search to automatically discover high-quality sub-network initializations, providing better starting points for pretraining. Third, we apply knowledge distillation from larger teacher models to speed up training and improve generalization. Together, these components make SLM pretraining substantially more efficient: our best model, discovered using evolutionary search and initialized with LLM weights, matches the validation perplexity of a comparable Pythia SLM while requiring $\mathbf{5.16\times}$ and $\mathbf{1.26\times}$ fewer floating point operations for token budgets of 10B and 100B, respectively. We release all code publicly, offering a practical and reproducible path toward cost-efficient small language model development at scale.
@@ 1 Introduction
  • Comprehensive analysis. We provide the first systematic comparison of sub-network initialization under knowledge distillation versus standard cross-entropy training, showing the benefits of knowledge distillation over standard pretraining. Our study spans multiple student scales and investigates how teacher size influences effectiveness of distillation.
  • Reproducible framework. We release an open-source library [^1] https://anonymous.4open.science/r/whittle-iclr-71CD/ for extracting sub-networks from existing LLM checkpoints. Together with our empirical findings, this establishes practical guidelines for compute-optimal SLM pretraining across different scales.
@@ 2.3 SLM Pretraining and Distillation
$\mathcal{L}_{\text{D}}=\sum_{i}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p_{t_{i}}}^{(T)}\log\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p_{t_{i}}}^{(T)}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}p_{s_{i}}}^{(T)}},$ (3)
@@ 3 Whittle: A Library for SLM Pre-training and Distillation
Given a pretrained decoder-only LLM from litgpt, we first convert it into a whittle model to enable flexible sub-network extraction. To evaluate a sub-network using the whittle model, we dynamically activate only structured components of the LLM associated with that sub-network using the set_sub_network() API (Listing 4). It allows the user to explicitly set architectural parameters of the sub-network, such as embedding dimension, intermediate size, number of heads, layers, query groups, and head size, as well as indices for sampled neurons, layers, and heads. Importantly, it allows to vary the number of heads, head size, intermediate size, and query groups across layers. This function is a core utility in whittle, supporting downstream procedures such as search, pretraining, and distillation.
@@
The search() API (Listing 2, Appendix F) constructs a whittle super-network from a base HuggingFace model and facilitates automated sub-network selection. It supports evolutionary strategies as well as algorithms from syne-tune (Salinas et al., 2022) [^4] https://github.com/syne-tune/syne-tune, and performs constrained search across parameter bins via rejection sampling. Each candidate sub-network is instantiated through set_sub_network() and evaluated on a task-specific metric, such as perplexity, to guide the search process.
@@
The convert_subnet_to_litgpt_model() function (Listing 3, Appendix F) transforms a selected sub-network configuration into a standalone GPT model within the litgpt framework. Given a super-network and a dictionary specifying architectural configurations (e.g., embedding dimension and number of heads), this utility extracts the corresponding sub-network and instantiates it as an independent GPT model. The resulting model can then be employed for downstream tasks such as pretraining, fine-tuning, or distillation.
@@
The pretrain() function (Listing 1, Appendix F) enables pretraining of a sub-network initialized from a checkpointed GPT model. Given the model weights, a configuration file describing the sub-network architecture, and a target dataset, this utility restores the model and resumes training from the specified state.
@@ 4.1 Search Space Definitions
Table 1 summarizes the search spaces for Pythia-6.9B and Pythia-12B, listing the allowable values for each transformer dimension. In the coarse layer-wise and fine-grained layer-wise settings, $h$, $h_{s}$, and $d$ are sampled independently at each layer. The fine-grained spaces extend this further with neuron-level sampling within each dimension, as detailed in Section 2.1. The size of each of the search spaces is listed in Table 16 (Appendix H).
@@ 4.2 Evolutionary Search for Optimal SLMs
We use perplexity on wikitext (Merity et al., 2017) as selection metric, with mutation and crossover probabilities fixed at $0.2$. For Pythia-6.9B and -12B, we define three bins with parameter counts of 385M–426M (bin-1), 961M–1.06B (bin-2), and 2.64B–2.91B (bin-3), centered at 5% of Pythia-410M, Pythia-1B, and Pythia-2.8B, respectively [^7] The bins were computed based on the exact number of parameters in the Pythia models. Each evolutionary run proceeds for 100 epochs, with results in the fine-grained layer-wise setting reported at the final epoch before rejection sampling becomes infeasible due to the combinatorial growth of the search space.
@@ 4.3 Pretraining of SLMs
Figure 4 presents the pretraining results. We compare pretraining of the extracted best sub-network (Supernet-init) against two baselines: (i) Random-init, where the same architecture is trained with random initialization and a 10B-token budget, and (ii) the original Pythia model (center of the bin), also trained with random initialization and the same budget. Across parameter bins, initializing from the supernet yields consistent improvements in validation perplexity. Notably, the gains are most pronounced for bin-3, indicating that supernet initialization is particularly beneficial in higher-parameter regimes, where our model achieved the same validation perplexity with $\mathbf{5.16\times}$ fewer FLOP. Results for a token-budget of 100B can be found in Appendix J.
@@ 4.4 Distillation of SLMs
Figure 5 illustrates the effect of distillation. We find that distillation consistently improves perplexity in both bin-1 and bin-2, with the model in bin-2. We also report the performance of distilled models on downstream tasks in Table 14 in Appendix D.
@@ 5 Ablations
Finally, in Figure 8, we evaluate different search metrics. Specifically, we compare activation-based importance scores (as in Minitron (Muralidharan et al., 2024)) and weight-magnitude scores (Han et al., 2015) against directly optimizing for perplexity in our setup. We define the details of the importance score computation procedure, i.e. the metric guiding the search, in Appendix E. All searches are run with for 100 epochs. We find that perplexity-based search consistently achieves lower perplexity than proxy metrics, suggesting that importance and magnitude scores are less reliable indicators of sub-network quality.
@@ Appendix A Related Work
Appendix B Additional Experimental Details
B.1 Hyperparameter Configurations of Experiments
@@
B.2 Computational Cost of the Evolutionary Search
In this section, we provide an overview of the cost overhead introduced by evolutionary search. In each bin, for every search space, we sample and evaluate a total of 5,050 subnetworks during the evolutionary search. For each candidate model, we computed the perplexity on 1,000 sequences of length 512. We approximate the average FLOP of the models in bins 1, 2 and 3 as the FLOPs of Pythia-410M, Pythia-1B, and Pythia-2.8B, since these models serve as the center of the bins. The total computational cost of the search for each search space is reported in Table 9. For comparison, the cost of pretraining the best model found in each bin on 10B tokens is as presented in Table 10. As Tables 9 and 10 indicate, the search phase consumes only a small fraction of the overall pretraining budget.
Bin Search Cost (exaFLOP)
1 2.3
2 5.4
3 15.4
Table 9: Cost of Evolutionary Search for different bins
Bin Pretraining Cost (exaFLOP)
1 20.9
2 63.3
3 176.6
Table 10: Cost of pretraining the best models in different bins
Revised Cost Savings.
We include the cost of the evolutionary search when computing the total cost savings achieved by our method. The updated FLOP-savings factors are reported below in Table 11. Additionally, we include the FLOP-savings factor considering only the pretraining budget in Table 12.
Bin FLOP Savings Factor
[…]
Table 11: FLOP-saving factors for all bins compared to pretraining the corresponding Pythia architectures.
Bin FLOP Savings Factor (excluding search cost)
[…]
Table 12: FLOP-saving factors for all bins compared to pretraining the corresponding Pythia architectures (without considering the search cost).
Furthermore, in Figure 9, we present the validation perplexity across different parameter ranges, taking the search cost into account.
Figure 9: Validation perplexity across different parameter ranges (offset with search cost).
@@ Appendix D Additional Results
Table 14: Evaluation of sub-networks extracted from Pythia-6.9b for bin-0 and bin-1. Reported numbers are metrics as defined in Section 4.5 (%). We compare training with the cross entropy loss ( from-supernet) to training with knowledge distillation (from-supernet-distill) loss.
@@ Appendix E Details on Importance Scoring
Importance scoring aims at defining scores for each transformer dimension, neuron or architecture parameter based on activation or weight magnitude. In our case, for a sub-network, the corresponding importance score serves as the proxy to sub-network quality or performance metrics like perplexity. The higher the importance score of a sub-network, the better its quality.
We adopt the dimension-wise importance scoring proposed by Muralidharan et al. [2024], which uses the activation of a component as proxy for its importance. Given a batch as input ${\bm{X}}\in\mathbb{R}^{B\times T\times d_{model}}$ after applying the embedding layer ${\bm{W}}^{emb}$ we compute the following scores for each component, where $B$ corresponds to the batch dimension and $T$ corresponds to the sequence length dimension, and absabs corresponds to the absolute value function:
@@ Appendix F Whittle APIs
An overview of the Whittle library is shown in Figure 10. In addition to the core functionalities of our framework described in Section 3, we provide an API to compute various importance metrics across different sub-network dimensions.
@@
Below we present the details of our API design for setting subnetwork 4, pretrain 1, convert to litgpt 3, distillation 5, 2 search and importance metric computation 6.
Listing 1: API for pretraining.
@@
Listing 2: API for supernetwork search.
@@
Listing 3: API for converting a subnet into a LitGPT model.
@@
Listing 4: API for activating a sub-network
def set_sub_network(
sub_network_n_embd: int = 4, # Embedding dim
sub_network_intermediate_size: int | list[int] = 42 # MLP size
sub_network_num_heads: int | list[int] = 4, # No. of Attention heads
sub_network_n_layers: int = 4, # No. of Layers
sub_network_query_groups: int | list[int] = 2, # No. of Query groups
sub_network_head_size: int | list[int] = 4, # Head size
sampled_layer_indices: list[int] = [2,3,5,6], # Sampled layers
@@
Appendix G Studying Distillation Hyperparameters
We define our distillation loss function below.
$\mathcal{L}=\alpha\,\mathcal{L}_{\text{CE}}(\mathbf{y},\mathbf{s})+\beta\sum_{i\in\mathcal{K}}\text{KL}\!\left(\text{softmax}\!\left(\tfrac{z_{t}^{(i)}}{T}\right)\,\|\,\text{softmax}\!\left(\tfrac{z_{s}^{(i)}}{T}\right)\right).,$ (5)
This loss has four tunable hyperparameters:
  1. 1.
    $\alpha$: the weight of the cross-entropy loss, which minimizes the entropy with respect to the ground-truth logits.
  2. 2.
    $\beta$: the weight of the KL-divergence term between the teacher and student logits.
  3. 3.
    $T$: the temperature parameter, which controls the smoothing of the teacher and student logit distributions.
  4. 4.
    $\mathcal{K}$: the number of top logits (Top-$\mathcal{K}$) used when computing the KL-divergence. A smaller $\mathcal{K}$ corresponds to a simpler distribution, while a larger $\mathcal{K}$ yields a more informative distribution. The maximum $\mathcal{K}$ equals the full number of teacher logits.
We now perform a grid-sweep over different choices of $\alpha$, $\beta$, $T$ and $\mathcal{K}$. We define the set of choices for $\alpha$ as $[0.2,0.8,0]$, the corresponding choices for $\beta$, which corresponds to $1-\alpha$ as $[0.8,0.2,1]$, the choices for temperature $T$ as $[0.8,0.9]$ and the choices for $\mathcal{K}$ as [1024,2048,num_teacher_logits][1024,2048,num\_teacher\_logits]. Figures 18-20 present the perplexity curves aggregated for different hyperparameter values. In general, we observe that using only the distillation loss, i.e., setting $\alpha=0$, is not recommended. Furthermore, a higher temperature and using the full logit distribution (Top-K = 0) perform best on average. In Table 15, we also present the importance of each of the hyperparameter choices and find that the most important one is the value of $\alpha$, followed by $\mathcal{K}$ and finally the temperature $T$.
Figure 18: Bin 1 - distillation hyperparameter importance
Figure 19: Bin 2 - distillation hyperparameter importance
Figure 20: Bin 3 - distillation hyperparameter importance
hyperparameter importance (mean) importance (std)
$\alpha$ 0.986442 0.011966
$\mathcal{K}$ 0.005968 0.000469
$T$ 0.005000 0.000551
Table 15: Hyperparameter Importance
Appendix H Search Space Sizes
The sizes of the search spaces for evolutionary search, especially for fine-grained (both uniform and layer-wise) can grow exponentially. In Table 16, we show the maximum number of configurations per search space for both of our base Pythia models (6.9B and 12B). For equations and information on how the number of search space configurations is calculated, refer to Section 2.1, and Table 1 for the configurations.
Table 16: Search Space Sizes (NN) for Pythia-6.9B and Pythia-12B Architectures
Search Space Pythia-6.9B ($N$) Pythia-12B ($N$)
Coarse Uniform $4.33\times 10^{12}$ $9.51\times 10^{12}$
Coarse Layer-wise $1.65\times 10^{244}$ $2.33\times 10^{281}$
Fine-grained Uniform $10^{14284}$ $10^{17841}$
Fine-grained Layer-wise $1.58\times 10^{159984}$} $5.31\times 10^{224611}$}
Appendix I Reproducibility Statements
@@
Appendix J Scaling Behavior of Subnetwork Extraction Under Larger Budgets
To assess how the cost savings from subnetwork extraction scale with substantially larger pretraining budgets, we trained the best model from Bin 2 and Bin 3 for 100 billion tokens and compared it against a Pythia-1B and Pythia-2.8B model, respectively, trained for the same number of tokens from random initialization. We summarize our key findings below.
  • Sustained FLOP Savings at the 100B-Token Scale. Our extracted subnetwork for Bin-3 achieves the same validation performance while requiring $1.26\times$ fewer FLOPs (a reduction of approximately 21%). Although this reduction is smaller than the $5.16\times$ savings observed at the 10B-token scale, it nevertheless demonstrates that subnetwork extraction continues to provide meaningful computational benefits even when the training budget is increased by an order of magnitude. This indicates that the method is not confined to low-budget regimes and remains competitive at significantly larger compute settings.
  • Improved Final Validation Perplexity. In addition to being more compute-efficient, the extracted model also attains a lower final validation perplexity. The baseline Pythia-2.8B reaches a perplexity of 11.397, while our model achieves 11.204.
  • Strong Downstream Performance Advantages. The extracted model outperforms the Pythia models trained from scratch across downstream evaluations. In particular, on MMLU-cont, our model achieves gains of up to 12% over the strongest Pythia baseline (see Table 17).
Base Model Initialization #Params COPA OpenBookQA Lambada-OpenAI Winogrande Social IQA MMLU-cont. MMLU CommonsenseQA PIQA ARC-challenge ARC-easy HellaSwag BoolQ Avg-acc
Pythia-6.9B Supernet Init 1.04B 74.00 38.40 45.604 56.98 41.81 32.54 27.04 20.88 75.03 38.57 71.76 60.05 61.19 49.53
Pythia-1B Random Init 1.01B 71.00 35.40 36.13 53.35 41.91 26.93 25.20 19.74 73.50 35.92 69.36 55.69 52.75 45.91
Pythia-6.9B Supernet Init 2.91B 76.00 37.40 53.23 58.01 42.02 34.45 38.04 43.41 77.69 41.38 73.57 64.01 63.49 54.05
Pythia-2.8B Random Init 2.78B 71.00 40.40 43.39 58.64 42.02 27.18 25.74 21.37 76.50 42.75 74.24 64.85 60.58 49.23
Table 17: Evaluation of Pythia models trained for 100B tokens across multiple benchmarks. Reported numbers are metrics as defined in Section 4.5 (%).
In Figures 21-23, we present the trajectories for validation perplexity, validation loss and train loss, respectively, for Bin 2 and Bin 3 architectures and the Pythia-based models trained from scratch for 100B tokens.
Figure 21: Bin 2 and Bin 3 validation perplexity for 100B token budget
Figure 22: Bin 2 and Bin 3 validation loss for 100B token budget
Figure 23: Bin 2 and Bin 3 train loss for 100B token budget